The sum of Raul's and Jesse's ages is 15. The product of their ages is 44. If Jesse is younger, how old is Jesse? answer is 4; x + y = 15; x(15-x)=44; -x^2-15x-44=0;so, x^2+15+44=0. Did I set it up right? Its a quadratic equation age problem. Please, please help/Thank you so much. Mimi
Answer by jorel1380(3719)

x+y=15
xy=44
x(15-x)=44
-x2+15x-44=0
x2-15x+44=0
(x-11)(x-4)=0
x=11 or 4
Raul is 11; Jesse is 4..